Solve by the elimination method
5r - 8s = -3
8r + 5s = 102
I came up with (9,6) r being 9 and s being 6. Is that correct?

Answer by ewatrrr(24785)


Hi, Yes! You are exactly corect
Solving by the elimination method
5r - 8s = -3
8r + 5s = 102
25/8r - 5s = -3*5/8 |multiplying 1st EQ by 5/8
8r + 5s = 102
89/8 r = -15/8 + 102
89r = -15 + 816
89r = 801
r = 9 and s = (5r+3)8 = 48/8=6